Analyze and compare Costco, Wal-Mart, and Nordstrom The general merchandise retail industry has a number of segments represen.................
Answer #1

For Nordstrom

Average inventory = ( Beginning inventory + Ending inventory)/2

= (1,733+1,945)/2

= $1,839 million

Inventory Turnover = Cost of goods sold/Average inventory

= 9,168/1,839

= 5

b)

Days sales in inventory = 365/Inventory turnover

= 365/5

= 73 days

Kindly comment if you need further assistance. Thanks‼!